Question 1) The economic concept of cost differs from the everyday notion of a monetary payment. In economics, the cost of doing something is what you have to give up to be able to do it. In this sense, you can face a cost even if you do not have to pay anyone anything. Economists refer to this as an opportunity cost. Suppose that a farmer has land that can produce 20 bushels of corn per acre or 10 bushels of wheat per acre. She currently is producing 100 bushels of corn and 100 bushels of wheat.

What is the opportunity cost to the farmer, measured in bushels of corn, of producing an additional 10 bushels of wheat (answer with numbers only. Ex: If the correct answer is 3 bushels then enter 3 in the box below)?
Answer #1

As the farmer's land can produce 20 bushels of corn or 10 bushels of wheat

the opportunity cost of producing 1 bushel of corn is 10 / 20 = 0.5 bushels of wheat.

Similarly, the opportunity cost for the farmer to grow an additional unit of wheat is 20 / 10 = 2 bushels of corn.

So, the opportunity cost to the farmer of producing an additional 10 bushels of wheat = 2*10 = 20 bushels of corn.

It means if the farmer wants to grow an additional 10 bushels of wheat i.e. 110 bushels of wheat, its corn production will be reduced to 100-20 = 80